A chemistry graduate student is given 125.mL of a 1.20M propanoic acid HC2H5CO2 solution. Propanoic acid is a weak acid with =Ka×1.310−5 What mass of KC2H5CO2 should the student dissolves in the HC2H5CO2 solution to turn it into a buffer with pH =5.13

Be sure your answer has a unit symbol, and round it to 2 significant digits.

?
Answer #1

Ka = 1.3*10^-5

pKa = - log (Ka)

= - log(1.3*10^-5)

= 4.886

use formula for buffer

pH = pKa + log ([KC2H5CO2]/[HC2H5CO2])

5.13 = 4.8861 + log ([KC2H5CO2]/[HC2H5CO2])

log ([KC2H5CO2]/[HC2H5CO2]) = 0.2439

[KC2H5CO2]/1.2 = 1.7537

[KC2H5CO2] = 2.1044

volume , V = 1.25*10^2 mL

= 0.125 L

use:

number of mol,

n = Molarity * Volume

= 2.104*0.125

= 0.263 mol

Molar mass of KC2H5CO2,

MM = 1*MM(K) + 3*MM(C) + 5*MM(H) + 2*MM(O)

= 1*39.1 + 3*12.01 + 5*1.008 + 2*16.0

= 112.17 g/mol

use:

mass of KC2H5CO2,

m = number of mol * molar mass

= 0.263 mol * 1.122*10^2 g/mol

= 29.51 g

Answer: 30. g